OpenStudy (anonymous):

use the substitution method to solve system of equations 2x+4y=14 x=3

OpenStudy (anonymous):

@geo12345679789

OpenStudy (anonymous):

thats not one of the answer choices. and where'd the 10 come from ?

OpenStudy (anonymous):

x=3 2x+4y=14 Now substitute x by 3: 2*3+4y=14 6+4y=14 4y=8 y=2

OpenStudy (anonymous):

@geo12345679789 I get the answer.

OpenStudy (anonymous):

I made a mistake, sorry. Now I have corrected it.

OpenStudy (anonymous):

thanks and its ok c:

OpenStudy (anonymous):

You are welcome!

OpenStudy (anonymous):

wait.. where'd you get the 8

OpenStudy (anonymous):

6+4y=14 Then 6+4y-6=14-6 So 4y=8

OpenStudy (anonymous):

@geo12345679789

OpenStudy (anonymous):

ah gotcha

OpenStudy (anonymous):

: )